A modern riffle can fire a bullet from 180 m/s to 1500 m/s.

a. Compute for the maximum height of a bullet fired by a riffle with an average velocity of 500m/s at an angle of 800.
1
Expert's answer
2020-12-14T07:28:55-0500

Let's write the equation of motion of the bullet in vertical direction:


"y=v_0tsin\theta-\dfrac{1}{2}gt^2, (1)"

here, "v_0 = 500\ \dfrac{m}{s}" is the initial velocity of the bullet, "t" is the time of flight of the bullet, "\theta=80^{\circ}" is the launch angle, "y" is the vertical displacement of the bullet (or the height) and "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's first find the time that the bullet takes to reach the maximum height from the kinematic equation:


"v_y=v_0sin\theta-gt_{rise},""0=v_0sin\theta-gt_{rise},""t_{rise}=\dfrac{v_0sin\theta}{g}."

Then, we can substitute "t_{rise}" into the first equation and find the maximum height:


"y_{max}=v_0sin\theta\cdot\dfrac{v_0sin\theta}{g}-\dfrac{1}{2}g(\dfrac{v_0sin\theta}{g})^2,""y_{max}=\dfrac{v_0^2sin^2\theta}{2g},""y_{max}=\dfrac{(500\ \dfrac{m}{s})^2\cdot sin^280^{\circ}}{2\cdot 9.8\ \dfrac{m}{s^2}}=12370\ m=12.37\ km."

Answer:

"y_{max}=12370\ m=12.37\ km."